OpenStudy (anonymous):

Need some help with my Functions & Linear Equations unit...I know how to graph coordinates and domain and range but now I have questions with dependent variables and replacement sets. I'm behind in my math, but I'm not gonna post all the questions and cheat I just want someone to maybe give me some quick definitions please?

OpenStudy (lucaz):

what kind of problem/question you have trouble to solve?

OpenStudy (anonymous):

Well heres one of them, The x-intercept of 3x + 2y = 12 is (__, 0). I know how to find x on a line chart, but not so sure about this

OpenStudy (lucaz):

x-intercept is when the line crosses the x-axis, when y = 0, you want to find the value of x, you already know that y is zero, replace y by 0 and solve for x

OpenStudy (the_fizicx99):

To find the x intercept set y for 0 To find the y intercept set x for 0 So to find x, 3x + 2(0) = 12 = 3x = 12 /3 / 3 x = 4 (4,0)

OpenStudy (lucaz):

pretty straight forward

OpenStudy (anonymous):

Okay, thank you!

OpenStudy (lucaz):

no problem